OpenStudy (anonymous):

Place the indicated product in the proper location on the grid. (a + 3)(a - 2)

OpenStudy (anonymous):

how do you mean proper location on the grid?

OpenStudy (anonymous):

nvmind just (a + 3)(a - 2) . ignore the grid part

OpenStudy (anonymous):

ok so you need to distribute using the foil method

OpenStudy (anonymous):

so (a + 3) ( a -2) = a*a + (-2*a) + 3*a + (3*-2)

OpenStudy (anonymous):

= a^2 -2a +3a -6 = a^2 +a - 6